OpenStudy (anonymous):

PLs. help me simplifying this: x^3+64y^3/(x^2-4xy+16y^2)(x^2-16y^2)

OpenStudy (anonymous):

holy moly

OpenStudy (anonymous):

my ans is.. \[1/x-4y\]

OpenStudy (anonymous):

i know im wrong

OpenStudy (anonymous):

numerator is the sum of two cubes so it factor as \[x^3+64y^3=(x+4y)(x^2-4xy+16)\]

OpenStudy (anonymous):

yah, i got that

OpenStudy (anonymous):

one factor cancels. then cancel the \[x+4y\] and you get \[\frac{1}{x-4y}\]

OpenStudy (anonymous):

ok looks like you got these no problem. factor and cancel gets it

OpenStudy (anonymous):

you are not wrong. you are right. done

OpenStudy (anonymous):

thank you again